Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

Given that P(A) = 0.3, P(B) = 0.6, and A and B are independent events, compute each probability. (b) P(A or B)

OpenStudy (anonymous):

P(A)+P(B)-P(A)x P(B) = 0.3+0.6-0.3 x 0.6 = 0.72

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!